OpenStudy (anonymous):

14-4g=6 solve the two step equations

OpenStudy (anonymous):

subtract 14 from both sides, then divide both sides by -4

OpenStudy (anonymous):

g=-2?

OpenStudy (anonymous):

\[g=\frac{6-14}{-4}\]

OpenStudy (anonymous):

g=2. 14-4g=6 I would add 4g to each side first, 14=6+4g then subtract 6 8=4g Now, divide by 4 to find g. 2=g Test the answer by inserting 2 for g in the expression. 14-4(2)=6 14-8=6 6=6

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Nancy.. g=2.. smh..

OpenStudy (anonymous):

sorry 2 not -2

OpenStudy (anonymous):

g=2

OpenStudy (anonymous):

i keep forget the sign